OpenStudy (wade123):

Write the sum using summation notation, assuming the suggested pattern continues. 8 + 27 + 64 + 125 + ... + n^3

OpenStudy (wade123):

my answer-

ganeshie8 (ganeshie8):

plugin n = 2, do you get the first term, 8 ?

OpenStudy (wade123):

no..

ganeshie8 (ganeshie8):

\(\large (n-1)^3\) plugin \(n=2\), what do u get ?

OpenStudy (wade123):

1

OpenStudy (wade123):

so it should be n^3

OpenStudy (wade123):

where n=2

ganeshie8 (ganeshie8):

So \(\large \sum \limits_{n=2}^{\infty} (n-1)^3 \) is not a correct representation for the given series

ganeshie8 (ganeshie8):

what are the other options ?

OpenStudy (wade123):

it is written

OpenStudy (wade123):

sorry my internet is a little slow

OpenStudy (wade123):

wouldnt it be this? @ganeshie8

OpenStudy (anonymous):

\[8+27+64+....+n^3=2^3+3^3+4^3+....+n^3\] \[=1^3+2^3+3^3+...+n^3-1^3=\sum n^3-1=\left( \sum n \right)^2-1\] \[=\left( \frac{ n \left( n+1 \right) }{ 2 } \right)^2-1\]

OpenStudy (wade123):

i sent that through and it was wrong..

OpenStudy (wade123):

why wouldnt mine be right?

ganeshie8 (ganeshie8):

thats the correct sum notation, but the upper limit has a mistake.

ganeshie8 (ganeshie8):

\(\large \sum \limits_{n=2}^{\color{red}{\infty}} n^3 \)

ganeshie8 (ganeshie8):

Notice that the given sequence is FINITE, so upper limit has to be "n"

OpenStudy (wade123):

yeah that makes sense

OpenStudy (wade123):

thank you so much

ganeshie8 (ganeshie8):

Here is the correct sum notation : \(\large \sum \limits_{i=2}^{\color{red}{n}} i^3 \)

OpenStudy (wade123):

if i get the question wrong multiple times, it gives me options and this is what it gave me. I am gonna go with A @ganeshie8

ganeshie8 (ganeshie8):

yes \(a\) is the only best option among the given options :)

OpenStudy (wade123):

yay (: